OpenStudy (anonymous):

You have 16 coins that are heads up and 18 coins that are tails up. After you add some coins that are tails up, the ratio of heads up to tails up is 1 to 1.5 How many coins are tails up? How many coins in all?

OpenStudy (anonymous):

\[16/( 18+x) = 1/1.5\]

OpenStudy (anonymous):

Then 18+ x = 1.5(16)

OpenStudy (anonymous):

18 + x = 24 x = 6 40 total coins

OpenStudy (anonymous):

24 Tails

OpenStudy (anonymous):

Thanks, Bob!

OpenStudy (anonymous):

First post -- sorry it wasn't "prettier ;-)